pfpcoh (cohomology/homology groups for

p F q) マニュアル

Edition : auto generated by oxgentexi on 30 March 2017

OpenXM.org

1 pFq に関する (コ) ホモロジー

この節では超幾何関数 pFq (_{p}F_{q}) の (コ) ホモロジ群に関連した不変量を計算する関数を解説する.

OpenXM/Risa/Asir での利用にあたっては,

load("pfpcoh.rr")$ load("pfphom.rr")$

が始めに必要.

1.0.1 pfp_omega
pfp_omega(P)

: It returns the Gauss-Manin connection Omega for the generalized hypergeometric function P F P-1 (aa1,aa2, ...; cc1, cc2, ...;x) .

Description:

Define a vector valued function Y of which elements are generalized hypergeometric function f 1=F and f 2=xdf 1/dx, f3=xd f 2/dx, ... It satisfies dY/dx= Omega Y. Generalized hypergeometric function is defined by the series p F p-1(aa1,aa2, ...; cc1, cc2, ...;x) = sum(k=0,infty; (aa1) k (aa2) k .../( (1) k (cc1) k (cc2) k ... ) x^k)

Example:

 pfp_omega(3);
1.0.2 pfpcoh_intersection
pfpcoh_intersection(P)

: pfpcoh intersection(P) returns an intersection matrix for cocycles associated to the generalized hypergeometric function p F (p-1).

Description:

This program pfpcoh.rr computes an intersection matrix S of cocycles of p F p-1 and compares it with the matrix obtained by solving a differential equation for intersection matrix.

Algorithm:

Ohara, Sugiki, Takayama, Quadratic Relations for Generalized Hypergeometric Functions p F p-1

Example:

load("pfpcoh.rr")$
S=pfpcoh_intersection(3);

Author : K.Ohara

1.0.3 pfphom_intersection
pfphom_intersection(P)

: intersection matrix of homology cycles.

Description:

Computing intersection matrix of cycles associated to p F (p-1). As to the meaning of parameters c1, c2, c3, ..., see the paper Ohara, Kyushu J. Math. Vol. 51 PP.123.

Algorithm:

Ohara, Sugiki, Takayama, Quadratic Relations for Generalized Hypergeometric Functions p F p-1

Example:

           SS = pfphom_intersection(3)$

You get the intersection matrix of homologies for 3 F 2. Author : K.Ohara

1.0.4 pfphom_monodromy_pair_kyushu
pfphom_monodromy_pair_kyushu(P)

: Description:

It returns the pair of monodromy matrices. Algorithm:

Ohara, Kyushu J. Math. Vol.51 PP.123 (1997) Example:

       MP = pfphom_monodromy_pair_kyushu(3)$

You get a pair of monodromy matricies for 3F2 standing for two paths encircling 0 and Index 3

Index

(Index is nonexistent) (Index is nonexistent)

Short Contents

1 pFq に関する (コ) ホモロジー . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Table of Contents

1 pFq に関する (コ) ホモロジー . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.0.1 pfp_omega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.0.2 pfpcoh_intersection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.0.3 pfphom_intersection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.0.4 pfphom_monodromy_pair_kyushu . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3